If an equipment costs 800,000 and lasts 7 years, what should be the minimum annual (equal) cash inflow before it's worthwhile to purchase the equipment? Assume that the cost of capital is 12%.
Answer #1

Let the annual inflow be C.

Initial investment, C0 = 800,000

N, number of periods involved = 7

Cost of capital, R = 12%

Hence, NPV of the equipment = - C0 + PV of annuity of C = - C0 + C / R x [1 - (1 + R)-N] = - 800,000 + C/12% x [1 - (1 + 12%)-7] = - 800,000 + 4.5638C

It's worthwhile to purchase the equipment if NPV ≥ 0

- 800,000 + 4.5638C ≥ 0

Hence, C ≥ 800,000 / 4.5638 = $  175,294

Hence, minimum annual equal inflow required to make the equipment worthwhile to purchase = Cmin = $ 175,294